Operating Leverage Beck Inc. and Bryant Inc have the following operating data: Beck Inc Bryant Inc. Sales Varlable costs Contribution margin Fixed costs Income from operations a. Compute the operating leverage for Beck Inc. and Bryant Inc. If required, round to one decimal place. Beck Inc. $1,250,000$2.000,000 1,250,000 $500,000$750,000 750,000 400,000 $100,000 $300,000 < ) Bryant Inc. 2.5 b. How much would income from operations increase for sach company if the sales of each increased by 209%7 It required, rounid answers to nearest whole number. Dollars Percentage Beck Inc 34 | X % Bryant Inc operating leverage meana c. The difference in the increases of income from operations is due to the difference in the operating leverages. Beck tnc,s higher that its fixed costs are a largerpercentage of contribution margin than are Brvant Inc-s Previous Neot Check My Wark
Answer #1

a. Operating leverage = Contribution margin / Income from operations

Beck Inc.

[ $500,000 / $100,000 ]

5

Bryant Inc.

[ $750,000 / $300,000 ]

2.5

b. % change in income from operations / % change in sales = Operating leverage

In case of both the companies sales is increased by 20%

Beck Inc. :

% change in income from operation / 20% = 5

% change in income from operations = 5 * 20% = 100%

Thus here if the sales is increased by 20%, the income from operations will increase by 100%

Increase in income from operations = Current income from operations * % increase in income from operations = $100,000 * 100% = $100,000

Bryant Inc. :

% change in income from operations / 20% = 2.5

% change in income from operations = 2.5 * 20% = 50%

Thus here if the sales is increased by 20%, the income from operations will increase by 50%

Increase in income from operations = Current income from operations * % increase in income from operations = $300,000 * 50% = $150,000

Dollars Percentage
Beck Inc. $100,000 100%
Bryant Inc. $150,000 50%

c. The difference in the increases in income from operations is due to the difference in the operating leverages. Beck Inc's higher operating leverage means that its fixed costs are a larger percentage of contribution margin than are Bryant Inc.'s